# SymMaP: Improving Computational Efficiency in Linear Solvers through Symbolic Preconditioning

### v1

@@ Abstract
Matrix preconditioning is a critical technique to accelerate the solution of linear systems, where performance heavily depends on the selection of preconditioning parameters. Traditional parameter selection approaches often define fixed constants for specific scenarios. However, they rely on domain expertise and fail to consider the instance-wise features for individual problems, limiting their performance. In contrast, machine learning (ML) approaches, though promising, are hindered by high inference costs and limited interpretability. To combine the strengths of both approaches, we propose a symbolic discovery framework—namely, Symbolic Matrix Preconditioning (SymMaP)—to learn efficient symbolic expressions for preconditioning parameters. Specifically, we employ a neural network to search the high-dimensional discrete space for expressions that can accurately predict the optimal parameters. The learned expression allows for high inference efficiency and excellent interpretability (expressed in concise symbolic formulas), making it simple and reliable for deployment. Experimental results show that SymMaP consistently outperforms traditional strategies across various benchmarks  [^1] Our code is available at https://github.com/Minghaom2/SymMaP..
@@ 1 Introduction
Linear systems are foundational in machine learning, physics, engineering, and other scientific fields (leon2006linear, ; leveque2007finite, ; DBLP:journals/jcphy/GaoW23, ; gao2025CROP, ). Since analytical solutions are often unavailable, efficient numerical algorithms become essential (demmel1997applied, ). Matrix preconditioning, a key technique in this domain, accelerates iterative solvers and improves computational stability (trefethen2022numerical, ; chen2005matrix, ). For instance, the successive over-relaxation (SOR) method optimizes convergence by integrating Gauss-Seidel iterations with a weighted update scheme governed by the over-relaxation factor $\omega$ (golub2013matrix, ).
@@ 2.1 Matrix Preconditioning Technique
Matrix preconditioning is a technique employed to accelerate the convergence of iterative solvers and enhance the stability of algorithms. It is generally employed in solving linear systems (chen2005matrix, ; golub2013matrix, ), which are typically expressed in the form: ${\bm{Ax}}=\bm{b}.$ The fundamental idea of preconditioning is to transform the original problem into an equivalent one with better numerical properties. Specifically, this technique involves finding a preconditioner $\bm{M}$ that approximates either the inverse of $\bm{A}$ or some form conducive to iterative solutions (chen2005matrix, ). Consequently, the original equation is transformed into $\bm{MAx}=\bm{Mb}.$ There are generally two optimization objectives: 1. To accelerate the convergence of iterations by altering the spectral distribution of the matrix $\bm{A}$. 2. To reduce the condition number of the matrix $\bm{A}$, thereby lessening its ill-conditioning and enhancing the stability of the iteration. Some common preconditioning techniques include the Jacobi, Gauss-Seidel, SOR (young1954iterative, ), algebraic multigrid (AMG) (trottenberg2000multigrid, ), etc.
@@ 2.2 Prefix Notation
In this notation, operators can be unary (e.g., $\sin$, $\cos$) or binary (e.g., ++, -, $\times$, $\div$), while operands can be constants or variables (landajuela2021discovering, ). Each prefix expression uniquely corresponds to a symbolic tree structure, facilitating the conversion back to the original mathematical expression (lample2019deep, ).
@@ 3 Motivation
The selection of matrix preconditioning parameters significantly affects their effectiveness (chen2005matrix, ). To design appropriate algorithmic prediction parameters, we first analyze the optimization space for preconditioning parameter selection and investigate the existence of optimal parameters. Next, we analyze the unique challenges present in this scenario. Finally, to address these challenges, we design a symbolic discovery framework to select the parameters.
@@ 3.1 Motivation for Optimizing Preconditioning Parameters
As illustrated in Figure 1, the choice of relaxation factors $\omega$ significantly impacts the iteration count and computation time, when solving a second-order elliptic partial differential equation (PDE) (evans2022partial, ) using SOR preconditioning (golub2013matrix, ). There exists an optimal parameter $\omega_{op}$ that minimizes the computation time, with specific details available in the Appendix B.2.
To further analyze the optimization space of preconditioning parameters, we evaluate the impact of various parameter selection strategies on preconditioning performance. As shown in Figure 1, the ’Optimal Parameter $\omega_{op}$’ represents the parameter that minimizes computation time in each experiment, serving as the theoretical upper limit of our optimization. The ’Optimal Fixed Constant’ refers to a fixed constant that minimizes average computation time, and ’Default Parameter’ corresponds to the default setting of $\omega=1$ in the portable extensible toolkit for scientific computation (PETSc) (petsc-web-page, ). The gap between the optimal fixed constant and the optimal parameter highlights significant potential for optimizing preconditioning parameter selection, motivating this paper. The performance of our SymMaP algorithm approaches the optimal parameter, demonstrating its accuracy in learning the optimal parameter expression.
@@ 3.2 Challenges in Predicting Efficient Preconditioning Parameters
(C1) Strong Generalization Capability: Real-world scientific computing scenarios vary significantly. For instance, the choice of PDE grid form can lead to significant variations in matrix structure (johnson2009numerical, ), resulting in distinct optimal parameters. Moreover, preconditioning addresses various optimization goals, such as reducing computational time, the number of iterations, and lowering condition numbers (chen2005matrix, ). This necessitates that parameter prediction algorithms possess robust generalization capabilities: they should take problem scenarios and features as inputs while applying them to different preconditioning methods and optimization goals.
(C2) Computational Efficiency: linear system solver typically relies on Krylov subspace methods implemented in low-level libraries optimized for CPU architectures, such as PETSc (petsc-web-page, ), LAPACK (laug, ). Algorithms like generalized minimal residual method (GMRES) (saad1986gmres, ) and conjugate gradient (CG) (greenbaum1997iterative, ) iteratively compute the matrix’s invariant subspace, favoring single-threaded or limited multi-threaded execution modes. Preconditioning techniques aim to accelerate these solvers without significant additional computational overhead, often adopting implicit iterative formats (e.g., SOR (chen2005matrix, )) or utilizing low-cost matrix decompositions (e.g., AMG (trottenberg2000multigrid, )). Therefore, any parameter prediction algorithms must be compatible with CPU environments and seamlessly integrated into existing algorithm libraries. At the same time, it must maintain low computational costs to preserve the performance benefits of preconditioning.
(C3) Algorithmic Transparency: Algorithms in scientific computing often require rigorous analysis under mathematical theories. Opaque prediction algorithms could confuse researchers. For instance, the relaxation factor $\omega$ in SOR needs to avoid being too close to 0 or 2 in some scenarios (agarwal2000difference, ). This is an issue that opaque algorithms cannot avoid in advance. Moreover, interpretable algorithms can guide researchers to conduct further studies and reveal the underlying mathematical structures of problems. Therefore, these pose challenges to the transparency and interpretability of the parameter prediction algorithms.
@@ 3.3 Symbolic Discovery to Preconditioning Parameter Selection
Firstly, symbolic discovery can accommodate various types of input parameters and can specifically tailor symbolic expression learning for different preconditioning methods and optimization goals (cranmer2020discovering, ), thereby meeting the requirement for broad applicability in scientific computing tasks (C1). Secondly, the explicit expressions derived are computationally lightweight and can be quickly evaluated at runtime. They integrate seamlessly into existing CPU-based algorithm libraries like PETSc (petsc-web-page, ) with almost no overhead (C2). Thirdly, the symbolic discovery provides transparent and interpretable expressions (rudin2019stop, ), allowing researchers to understand the influence of parameters within existing theoretical frameworks and identify potential numerical stability issues. This interpretability fosters trust in the algorithm’s predictions and supports further theoretical exploration (C3).
@@ 4.2 The Generation of Symbolic Expressions
Token Library. For SymMaP, we define the library $\mathcal{L}$ of mathematical operators and operands as $\{+,-,\times,\div,\text{sqrt},$ $\exp,\log,\text{pow},1.0\}$. Although other operators such as poly, $\sin$ and $\cos$ are frequently used (udrescu2020ai, ), we decided to exclude them because they offer limited explanatory power in matrix preconditioning and significantly increase the time and memory consumption during training.
After converting the mathematical expressions into prefix notation, we use this tokenized representation as a pre-order traversal of the expression tree (zaremba2014learning, ). In each iteration, the RNN receives a pair consisting of a parent node and a sibling node as inputs. Then the RNN outputs a categorical distribution over all possible next tokens. The parent node refers to the last incomplete operator that requires additional operands to form a complete expression. The sibling node, in the context of a binary operator, represents the operand that has already been processed and incorporated into the expression. In cases where no parent or sibling node is applicable, they are designated as empty nodes. This structured input method enables the RNN to maintain contextual awareness and effectively predict the sequence of tokens that form valid mathematical expressions.
The Sequential Model. During the generation of a single symbolic expression, the RNN emits a categorical distribution for each "next token" at each step. This distribution is represented as a vector $\psi^{(i)}_{\bm{{\theta}}}$, where $i$ denotes the $i$-th step and $\bm{{\theta}}$ represents the parameters of the RNN. The elements of the vector correspond to the probabilities of each token, conditioned on the previously selected tokens in the traversal (petersen2019deep, ):
@@
Here, $\bm{\tau}_{i}$ denotes the index of the token selected at the $i$-th step. The probability of generating the entire symbolic expression $\bm{\tau}$ is then the product of the conditional probabilities of all tokens (petersen2019deep, ; landajuela2021discovering, ):
@@ 4.3 The Reward Function
Once a symbolic expression is fully generated (i.e., the symbolic tree reaches all its leaf nodes), we evaluate its fitness by calculating the normalized root-mean-square error (NRMSE), a metric frequently used in genetic programming symbolic discovery (schmidt2009distilling, ). It is defined as $\text{NRMSE}=\frac{1}{\sigma_{y}}\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}$ , where $\hat{y}_{i}=\bm{\tau}(x_{i})$ is the predicted value for the $i$-th sample, $x_{i}$ is the problem feature parameter, $y_{i}$ is the optimal preconditioning parameter, $\sigma_{y}$ is the standard deviation of the target values $y$, and $n$ is the number of data. To bound this fitness value between 0 and 1, we apply a squashing function: $R(\bm{\tau})=1/(1+\text{NRMSE}).$ Our objective is to maximize $R(\bm{\tau})$, thereby minimizing the NRMSE and improving the accuracy of the generated expressions.
@@ 4.4 The Training Algorithm
Risk-seeking Policy. It is natural to consider the expectation of the reward function, i.e., $\operatorname{\mathbb{E}}_{\bm{\tau}\sim\bm{p}(\bm{\tau};\bm{\theta})}[R(\bm{\tau})]$, as the objective function to optimize. We can apply the ’log-integral’ trick (williams1992simple, ) and obtain
@@
In the context of symbolic regression, model performance is often driven by a few exceptional results that outperform others by a significant margin (petersen2019deep, ; tamar2015policy, ). With this in mind, we adopt a risk-seeking policy, which aims to maximize:
@@
where $\text{CDF}(R(\bm{\tau});\bm{\theta})$ refers to the cumulative distribution function. From this, the gradient of $J(\bm{\theta},\varepsilon)$ can be derived as (petersen2019deep, ):
@@ 5 Experiments
Preconditioners: We considered three different preconditioners and various optimization metrics: 1. SOR preconditioner with the relaxation factor $\omega$ (golub2013matrix, ); 2. SSOR preconditioner with the relaxation factor $\omega$ (golub2013matrix, ); 3. AMG preconditioner with the threshold parameters $\theta_{T}$ (trottenberg2000multigrid, ).
Datasets: We investigated linear systems derived from five distinct PDE classes: 1. Darcy Flow Problems (li2020fourier, ), 2. Second-order Elliptic PDEs (evans2022partial, ), 3. Biharmonic Equations (barrata2023dolfinx, ), 4. Thermal Problems (wang2024accelerating, ). 5. Poisson Equations (wang2024accelerating, ). All cases except biharmonic equations yield symmetric matrices. Notably, the non-symmetric matrices from biharmonic equations are incompatible with SSOR and AMG preconditioning techniques. Additionally, we conducted partial tests on the Lippmann-Schwinger equation (a numerical integration problem) and Markov chains (an optimization problem). Details of these experiments can be found in Appendix E.7.
Baselines: We compared SymMaP against various parameter selection methods for preconditioning. Specifically, the comparison involved the following scenarios: 1. No matrix preconditioning, 2. Default parameters in PETSc (petsc-web-page, ), 3. Fixed constants, 4. Optimized fixed constants.
Experiment Settings: All preconditioning procedures were uniformly implemented using the C-based PETSc library (petsc-web-page, ) to maintain evaluation consistency. The experiments were conducted within PETSc’s linear solver framework (GMRES and CG) (golub2013matrix, ), with condition numbers computed through the built-in function KSPComputeExtremeSingularValues.
@@ A.1 Machine Learning for Algorithm Discovery
Machine learning has the potential to uncover implicit rules beyond human intuition from training data, enabling the construction of algorithms that outperform handcrafted programs. Approaches to algorithm discovery in machine learning encompass symbolic discovery, program search, and more. Specifically, program search focuses on optimizing the computational processes of algorithms. For example, [44] explores the discovery of faster sorting algorithms, while [8] investigates efficient optimization algorithms.
In contrast, symbolic discovery aims to search within the space of small mathematical expressions rather than computational streams [46, 30]. This approach is analogous to an extreme form of model distillation, where knowledge extracted from black-box neural networks is distilled into explicit mathematical expressions. Traditional methods for symbolic discovery have relied on evolutionary algorithms, including genetic programming [47]. Recently, deep learning has emerged as a powerful tool in this domain, offering enhanced representational capacity and new avenues for solving symbolic discovery problems [54, 10].

### v2

@@ Abstract
Matrix preconditioning is a critical technique to accelerate the solution of linear systems, where performance heavily depends on the selection of preconditioning parameters. Traditional parameter selection approaches often define fixed constants for specific scenarios. However, they rely on domain expertise and fail to consider the instance-wise features for individual problems, limiting their performance. In contrast, machine learning (ML) approaches, though promising, are hindered by high inference costs and limited interpretability. To combine the strengths of both approaches, we propose a symbolic discovery framework-namely, Symbolic Matrix Preconditioning (SymMaP)-to learn efficient symbolic expressions for preconditioning parameters. Specifically, we employ a neural network to search the high-dimensional discrete space for expressions that can accurately predict the optimal parameters. The learned expression allows for high inference efficiency and excellent interpretability (expressed in concise symbolic formulas), making it simple and reliable for deployment. Experimental results show that SymMaP consistently outperforms traditional strategies across various benchmarks  [^1] Our code is available at https://github.com/Minghaom2/SymMaP..
@@ 1 Introduction
Linear systems are foundational in machine learning, physics, engineering, and other scientific fields (leon2006linear; leveque2007finite; DBLP:journals/jcphy/GaoW23; gao2025CROP). Since analytical solutions are often unavailable, efficient numerical algorithms become essential (demmel1997applied). Matrix preconditioning, a key technique in this domain, accelerates iterative solvers and improves computational stability (trefethen2022numerical; chen2005matrix). For instance, the successive over-relaxation (SOR) method optimizes convergence by integrating Gauss-Seidel iterations with a weighted update scheme governed by the over-relaxation factor $\omega$ (golub2013matrix).
@@ 2.1 Matrix Preconditioning Technique
Matrix preconditioning is a technique employed to accelerate the convergence of iterative solvers and enhance the stability of algorithms. It is generally employed in solving linear systems (chen2005matrix; golub2013matrix), which are typically expressed in the form: ${\bm{Ax}}=\bm{b}.$ The fundamental idea of preconditioning is to transform the original problem into an equivalent one with better numerical properties. Specifically, this technique involves finding a preconditioner $\bm{M}$ that approximates either the inverse of $\bm{A}$ or some form conducive to iterative solutions (chen2005matrix). Consequently, the original equation is transformed into $\bm{MAx}=\bm{Mb}.$ There are generally two optimization objectives: 1. To accelerate the convergence of iterations by altering the spectral distribution of the matrix $\bm{A}$. 2. To reduce the condition number of the matrix $\bm{A}$, thereby lessening its ill-conditioning and enhancing the stability of the iteration. Some common preconditioning techniques include the Jacobi, Gauss-Seidel, SOR (young1954iterative), algebraic multigrid (AMG) (trottenberg2000multigrid), etc.
@@ 2.2 Prefix Notation
In this notation, operators can be unary (e.g., $\sin$, $\cos$) or binary (e.g., ++, -, $\times$, $\div$), while operands can be constants or variables (landajuela2021discovering). Each prefix expression uniquely corresponds to a symbolic tree structure, facilitating the conversion back to the original mathematical expression (lample2019deep).
@@ 3 Motivation
The selection of matrix preconditioning parameters significantly affects their effectiveness (chen2005matrix). To design appropriate algorithmic prediction parameters, we first analyze the optimization space for preconditioning parameter selection and investigate the existence of optimal parameters. Next, we analyze the unique challenges present in this scenario. Finally, to address these challenges, we design a symbolic discovery framework to select the parameters.
@@ 3.1 Motivation for Optimizing Preconditioning Parameters
As illustrated in Figure 1, the choice of relaxation factors $\omega$ significantly impacts the iteration count and computation time, when solving a second-order elliptic partial differential equation (PDE) (evans2022partial) using SOR preconditioning (golub2013matrix). There exists an optimal parameter $\omega_{op}$ that minimizes the computation time, with specific details available in the Appendix B.2.
To further analyze the optimization space of preconditioning parameters, we evaluate the impact of various parameter selection strategies on preconditioning performance. As shown in Figure 1, the ’Optimal Parameter $\omega_{op}$’ represents the parameter that minimizes computation time in each experiment, serving as the theoretical upper limit of our optimization. The ’Optimal Fixed Constant’ refers to a fixed constant that minimizes average computation time, and ’Default Parameter’ corresponds to the default setting of $\omega=1$ in the portable extensible toolkit for scientific computation (PETSc) (petsc-web-page). The gap between the optimal fixed constant and the optimal parameter highlights significant potential for optimizing preconditioning parameter selection, motivating this paper. The performance of our SymMaP algorithm approaches the optimal parameter, demonstrating its accuracy in learning the optimal parameter expression.
@@ 3.2 Challenges in Predicting Efficient Preconditioning Parameters
(C1) Strong Generalization Capability: Real-world scientific computing scenarios vary significantly. For instance, the choice of PDE grid form can lead to significant variations in matrix structure (johnson2009numerical), resulting in distinct optimal parameters. Moreover, preconditioning addresses various optimization goals, such as reducing computational time, the number of iterations, and lowering condition numbers (chen2005matrix). This necessitates that parameter prediction algorithms possess robust generalization capabilities: they should take problem scenarios and features as inputs while applying them to different preconditioning methods and optimization goals.
(C2) Computational Efficiency: linear system solver typically relies on Krylov subspace methods implemented in low-level libraries optimized for CPU architectures, such as PETSc (petsc-web-page), LAPACK (laug). Algorithms like generalized minimal residual method (GMRES) (saad1986gmres) and conjugate gradient (CG) (greenbaum1997iterative) iteratively compute the matrix’s invariant subspace, favoring single-threaded or limited multi-threaded execution modes. Preconditioning techniques aim to accelerate these solvers without significant additional computational overhead, often adopting implicit iterative formats (e.g., SOR (chen2005matrix)) or utilizing low-cost matrix decompositions (e.g., AMG (trottenberg2000multigrid)). Therefore, any parameter prediction algorithms must be compatible with CPU environments and seamlessly integrated into existing algorithm libraries. At the same time, it must maintain low computational costs to preserve the performance benefits of preconditioning.
(C3) Algorithmic Transparency: Algorithms in scientific computing often require rigorous analysis under mathematical theories. Opaque prediction algorithms could confuse researchers. For instance, the relaxation factor $\omega$ in SOR needs to avoid being too close to 0 or 2 in some scenarios (agarwal2000difference). This is an issue that opaque algorithms cannot avoid in advance. Moreover, interpretable algorithms can guide researchers to conduct further studies and reveal the underlying mathematical structures of problems. Therefore, these pose challenges to the transparency and interpretability of the parameter prediction algorithms.
@@ 3.3 Symbolic Discovery to Preconditioning Parameter Selection
Firstly, symbolic discovery can accommodate various types of input parameters and can specifically tailor symbolic expression learning for different preconditioning methods and optimization goals (cranmer2020discovering), thereby meeting the requirement for broad applicability in scientific computing tasks (C1). Secondly, the explicit expressions derived are computationally lightweight and can be quickly evaluated at runtime. They integrate seamlessly into existing CPU-based algorithm libraries like PETSc (petsc-web-page) with almost no overhead (C2). Thirdly, the symbolic discovery provides transparent and interpretable expressions (rudin2019stop), allowing researchers to understand the influence of parameters within existing theoretical frameworks and identify potential numerical stability issues. This interpretability fosters trust in the algorithm’s predictions and supports further theoretical exploration (C3).
@@ 4.2 The Generation of Symbolic Expressions
Token Library. For SymMaP, we define the library $\mathcal{L}$ of mathematical operators and operands as $\{+,-,\times,\div,\text{sqrt},$ $\exp,\log,\text{pow},1.0\}$. Although other operators such as poly, $\sin$ and $\cos$ are frequently used (udrescu2020ai), we decided to exclude them because they offer limited explanatory power in matrix preconditioning and significantly increase the time and memory consumption during training.
After converting the mathematical expressions into prefix notation, we use this tokenized representation as a pre-order traversal of the expression tree (zaremba2014learning). In each iteration, the RNN receives a pair consisting of a parent node and a sibling node as inputs. Then the RNN outputs a categorical distribution over all possible next tokens. The parent node refers to the last incomplete operator that requires additional operands to form a complete expression. The sibling node, in the context of a binary operator, represents the operand that has already been processed and incorporated into the expression. In cases where no parent or sibling node is applicable, they are designated as empty nodes. This structured input method enables the RNN to maintain contextual awareness and effectively predict the sequence of tokens that form valid mathematical expressions.
The Sequential Model. During the generation of a single symbolic expression, the RNN emits a categorical distribution for each "next token" at each step. This distribution is represented as a vector $\psi^{(i)}_{\bm{{\theta}}}$, where $i$ denotes the $i$-th step and $\bm{{\theta}}$ represents the parameters of the RNN. The elements of the vector correspond to the probabilities of each token, conditioned on the previously selected tokens in the traversal (petersen2019deep):
@@
Here, $\bm{\tau}_{i}$ denotes the index of the token selected at the $i$-th step. The probability of generating the entire symbolic expression $\bm{\tau}$ is then the product of the conditional probabilities of all tokens (petersen2019deep; landajuela2021discovering):
@@ 4.3 The Reward Function
Once a symbolic expression is fully generated (i.e., the symbolic tree reaches all its leaf nodes), we evaluate its fitness by calculating the normalized root-mean-square error (NRMSE), a metric frequently used in genetic programming symbolic discovery (schmidt2009distilling). It is defined as $\text{NRMSE}=\frac{1}{\sigma_{y}}\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}$ , where $\hat{y}_{i}=\bm{\tau}(x_{i})$ is the predicted value for the $i$-th sample, $x_{i}$ is the problem feature parameter, $y_{i}$ is the optimal preconditioning parameter, $\sigma_{y}$ is the standard deviation of the target values $y$, and $n$ is the number of data. To bound this fitness value between 0 and 1, we apply a squashing function: $R(\bm{\tau})=1/(1+\text{NRMSE}).$ Our objective is to maximize $R(\bm{\tau})$, thereby minimizing the NRMSE and improving the accuracy of the generated expressions.
@@ 4.4 The Training Algorithm
Risk-seeking Policy. It is natural to consider the expectation of the reward function, i.e., $\operatorname{\mathbb{E}}_{\bm{\tau}\sim\bm{p}(\bm{\tau};\bm{\theta})}[R(\bm{\tau})]$, as the objective function to optimize. We can apply the ’log-integral’ trick (williams1992simple) and obtain
@@
In the context of symbolic regression, model performance is often driven by a few exceptional results that outperform others by a significant margin (petersen2019deep; tamar2015policy). With this in mind, we adopt a risk-seeking policy, which aims to maximize:
@@
where $\text{CDF}(R(\bm{\tau});\bm{\theta})$ refers to the cumulative distribution function. From this, the gradient of $J(\bm{\theta},\varepsilon)$ can be derived as (petersen2019deep):
@@ 5 Experiments
Preconditioners: We considered three different preconditioners and various optimization metrics: 1. SOR preconditioner with the relaxation factor $\omega$ (golub2013matrix); 2. SSOR preconditioner with the relaxation factor $\omega$ (golub2013matrix); 3. AMG preconditioner with the threshold parameters $\theta_{T}$ (trottenberg2000multigrid).
Datasets: We investigated linear systems derived from five distinct PDE classes: 1. Darcy Flow Problems (li2020fourier), 2. Second-order Elliptic PDEs (evans2022partial), 3. Biharmonic Equations (barrata2023dolfinx), 4. Thermal Problems (wang2024accelerating). 5. Poisson Equations (wang2024accelerating). All cases except biharmonic equations yield symmetric matrices. Notably, the non-symmetric matrices from biharmonic equations are incompatible with SSOR and AMG preconditioning techniques. Additionally, we conducted partial tests on the Lippmann-Schwinger equation (a numerical integration problem) and Markov chains (an optimization problem). Details of these experiments can be found in Appendix E.7.
Baselines: We compared SymMaP against various parameter selection methods for preconditioning. Specifically, the comparison involved the following scenarios: 1. No matrix preconditioning, 2. Default parameters in PETSc (petsc-web-page), 3. Fixed constants, 4. Optimized fixed constants.
Experiment Settings: All preconditioning procedures were uniformly implemented using the C-based PETSc library (petsc-web-page) to maintain evaluation consistency. The experiments were conducted within PETSc’s linear solver framework (GMRES and CG) (golub2013matrix), with condition numbers computed through the built-in function KSPComputeExtremeSingularValues.
@@ A.1 Machine Learning for Algorithm Discovery
Machine learning has the potential to uncover implicit rules beyond human intuition from training data, enabling the construction of algorithms that outperform handcrafted programs. Approaches to algorithm discovery in machine learning encompass symbolic discovery, program search, and more. Specifically, program search focuses on optimizing the computational processes of algorithms. For example, [mankowitz2023faster] explores the discovery of faster sorting algorithms, while [chen2024symbolic] investigates efficient optimization algorithms.
In contrast, symbolic discovery aims to search within the space of small mathematical expressions rather than computational streams [petersen2019deep, landajuela2021discovering]. This approach is analogous to an extreme form of model distillation, where knowledge extracted from black-box neural networks is distilled into explicit mathematical expressions. Traditional methods for symbolic discovery have relied on evolutionary algorithms, including genetic programming [poli2008field]. Recently, deep learning has emerged as a powerful tool in this domain, offering enhanced representational capacity and new avenues for solving symbolic discovery problems [schmidt2009distilling, cranmer2020discovering].